To dilute 150 ml of 1 F CaCl2 solution to 0.02 F would require how many milliliters of water?


1
Expert's answer
2021-02-28T06:15:33-0500

CM = n/V

n1 = n2

CM1 x V1 = CM2 x V2

1 x 0.150 = 0.02 x V2

V2 = (1 x 0.150)/0.02 = 7.5 L = 7 500 ml

The required amount of water is: 7 500 - 150 = 7 350 ml